# CoTracker: It is Better to Track Together

### v2

@@ Abstract
We introduce CoTracker, a transformer-based model that tracks dense points in a frame jointly across a video sequence. This differs from most existing state-of-the-art approaches that track points independently, ignoring their correlation. We show that joint tracking results in a significantly higher tracking accuracy and robustness.
We also provide several technical innovations, including the concept of virtual tracks, which allows CoTracker to track 70k points jointly and simultaneously. Furthermore, CoTracker operates causally on short windows (hence, it is suitable for online tasks), but is trained by unrolling the windows across longer video sequences, which enables and significantly improves long-term tracking.
We demonstrate qualitatively impressive tracking results, where points can be tracked for a long time even when they are occluded or leave the field of view. Quantitatively, CoTracker outperforms all recent trackers on standard benchmarks, often by a substantial margin. Code and model weights are available at: https://co-tracker.github.io/
Figure 1: CoTracker is a new dense point tracker. It can track 70k points jointly on a single GPU and exploit the correlation between different tracks to improve its performance. It has exceptional long-term tracking performance even in the presence of occlusions or when points leave the camera view (examples from DAVIS [32], the bottom row shows object tracks only).
@@ 1 Introduction
Establishing point correspondences is a fundamental problem in computer vision, necessary for many downstream tasks. Here, we are interested in estimating point correspondences in videos containing dynamic objects and a moving camera. Specifically, given one or more 2D points which are the projection of certain physical points of the underlying 3D scene, the goal is to find the location of the same physical points in all other frames of the video.
@@
There are two variants of this problem. In optical flow, the objective is to estimate the velocity of all points within a video frame. This estimation is performed jointly for all points, but the motion is only predicted at an infinitesimal distance. In tracking, the goal is to estimate the motion of points over an extended period of time. For efficiency, tracking methods typically focus on a sparse set of points and treat them as statistically independent. Even recent techniques such as TAPIR [11] and PIPs++ [52], which employ modern architectures and can track points even in the presence of occlusions, model tracks independently. However, this approximation is crude as points are often strongly correlated (e.g., when they belong to the same object).
In this paper, we hypothesise that accounting for the correlation between tracked points can significantly improve performance. Thus, we introduce CoTracker, a new neural tracker that, compared to prior deep networks for optical flow and point tracking, supports joint estimation of a very large number of tracks simultaneously. CoTracker outperforms prior trackers on all common tracking benchmarks by substantial margins, is based on a simple and extensible architecture, and is as flexible as prior solutions for single-point tracking, meaning that it can track arbitrary points, selected at any spatial location and time in the video, and support arbitrary configurations of such points.
We show that tracking more points jointly significantly improves tracking metrics. In fact, we also show that injecting additional support points (i.e., points that are not requested by the user but help the tracker) can further improve performance, and we study several support configurations.
@@
CoTracker achieves its high performance also due to several other innovations in its design and learning formulation. The network is a transformer operating in a sliding window fashion on a two-dimensional representation of tokens, where dimensions are time and the set of tracked points. Via suitable self-attention operators, the transformer can consider each track as a whole for the duration of a window and can exchange information between tracks, thus exploiting their correlation. When tracking a very large number of points, furthermore, CoTracker introduces in its transformer design a small number of representative virtual track tokens, switching from expensive self-attention between virtual tracks to cross-attention between the latter and the real ones. This way, CoTracker can jointly track a near-dense set of tracks on a single GPU.
@@
Another important contribution is unrolled training. CoTracker is an online algorithm that processes videos in a serial fashion via a sliding window. Within a window, the tracks are initialized with queried points, and the network is tasked with progressively refining these initial estimations through iterative applications of the transformer. Each subsequent overlapping window starts with refined predictions of the previous window and updates tracks for the new frames. We mask out points that have not yet been initialized in the current sliding window, which allows initializing tracks from any point in the video, including in the middle of a window, and to join windows together. We show that, by training the network by unrolling several sliding windows, CoTracker achieves excellent long-term tracking performance as well.
We train CoTracker on the synthetic TAP-Vid-Kubric [10] and evaluate it on three different benchmarks, TAP-Vid-DAVIS, PointOdyssey [52] and DynamicReplica [23]. Our architecture works well for tracking single points and excels for groups of points, obtaining state-of-the-art tracking performance in several benchmarks.
@@ 2 Related work
Optical flow estimates dense instantaneous motion. Originally approached by studying the colour constancy equation [17, 27, 3, 4], optical flow is nowadays tackled by using deep learning, starting with FlowNet [12, 19]. More recently, DCFlow [48] has introduced the computation of a 4D cost volume, used in most follow-up works [41, 45, 43]. Notable is RAFT [43], which introduced incremental flow updates and inspired a wave of subsequent works [22, 47, 22, 49]. We also adopt 4D cost volumes and iterated updates, but apply them to tracking.
Transformers [44] have also been applied to the optical flow problem [18, 36, 50]. Flowformer [18] drew inspiration from RAFT and proposed a transformer-based approach that tokenizes the 4D cost volume. GMFlow [50] replaced the update network with a softmax with self-attention for refinement. Perceiver IO [20] proposed a unified transformer for several tasks including optical flow.
Optical flow can in principle be applied to tracking by integrating predictions over time, but this open-loop approach accumulates errors in the form of drift, which motivates developing dedicated architectures like ours.
@@
There have been several approaches [33, 35, 21, 28] to extend optical flow to multiple frames. Traditionally, Kalman filtering [9, 13] was a staple mechanism to ensure temporal consistency. Modern multi-frame methods produce dense flow. RAFT [43] can be applied in a warm-start manner [39, 42] for multi-frame optical flow estimation. However, these methods are not designed for long-term tracking and do not consider points occluded for a long time. VideoFlow [35] extends optical flow to three and five consecutive frames by integrating forward and backward motion features during flow refinement. MFT [31] conducts optical flow estimation between distant frames and chooses the most reliable chain of optical flows. Our tracker scales to produce semi-dense tracks.
Figure 2: CoTracker architecture. We compute convolutional features $\phi(I_{t})$ for every frame and process them with sliding windows. In order to initialize track features $Q$, we bilinearly sample from $\phi(I_{t})$ with starting point locations $P$. Locations $P$ also serve to initialize estimated tracks $\hat{P}$. See Fig. 3 for a detailed visualization of one sliding window.
Joint Tracking.
Before the emergence of deep learning, some authors proposed handcrafted joint trackers [2, 34], but there is less work on doing so with modern deep network architectures like ours. Our work is weakly related to multiple object tracking [8] where tracking through occlusions [51], with changes in appearance [29], and with temporal priors [37] have been extensively studied. However, our focus is the tracking of points, not objects.
@@
TAP-Vid [10] introduced the problem of tracking any physical point in a video and proposed a benchmark and a simple baseline method for it, but is unable to track occluded points. PIPs [15] revisits the classic Particle Video [34] problem by introducing a model for point tracking through occlusions. The method tracks selected points with a fixed sliding window and restarts from the last frame where the point is visible. However, the model loses the target if it stays occluded beyond the size of the window. While the original Particle Video does track points jointly, PIPs and TAP-Vid track points in parallel, but independently. TAPIR [11] proposes a feed-forward point tracker with a matching stage inspired by TAP-Vid [10] and a refinement stage inspired by PIPs [15], which further improves tracking. PointOdyssey [52] addresses long-term tracking with PIPs++, a simplified version of PIPs, and introduces a benchmark for long-term tracking. However, both of these methods still track points independently. OmniMotion [46] optimizes a volumetric representation for each video, refining estimated correspondences in a canonical space. However, this approach requires test-time optimization, which, due to its cost, is not suitable for many practical applications, especially online ones.
Trackers and optical flow estimators are often trained using synthetic datasets [24, 30, 12, 40, 5, 52], as annotating real data can be challenging. Synthetic datasets provide accurate annotations, and training on them has demonstrated the ability to generalize to real-world data [30, 40].
@@ 3 CoTracker
Figure 3: CoTracker architecture. Visualization of one sliding window with $M$ iterative updates. During one iteration, we update point tracks $\hat{P}^{(m)}$ and track features $Q^{(m)}$. $Q^{(0)}$ is initialized with the initially sampled features $Q$ for all sliding windows, $\hat{P}^{(0)}$ with the starting locations for the first window. For other windows, $\hat{P}^{(0)}$ starts with predictions for frames processed in the preceding sliding window, and with the last predicted positions for the unseen frames. We compute visibility $\hat{v}$ after the last update $M$.
Our goal is to track 2D points throughout the duration of a video $V=(I_{t})_{t=1}^{T}$, which is a sequence of $T$ RGB frames $I_{t}\in\mathbb{R}^{3\times H\times W}$. The goal of the tracker is to predict $N$ point tracks $P^{i}_{t}=(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}$, $t=t^{i},\dots,T$, $i=1,\dots,N$, where $t^{i}\in\{1,\dots,T\}$ is the time when the track starts. The tracker also predicts the visibility flag $v^{i}_{t}\in\{0,1\}$ which tells if a point is visible or occluded in a given frame. To make the task definition unambiguous [10], we assume each point is visible at the start of their track (i.e., $v^{i}_{t_{i}}=1$). The tracker is thus given as input the video $V$ and the starting locations and times $(P^{i}_{t^{i}},t^{i})_{i=1}^{N}$ of $N$ tracks, and outputs an estimate $(\hat{P}^{i}_{t}=(\hat{x}^{i}_{t},\hat{y}^{i}_{t}),\hat{v}^{i}_{t})$ of the track locations and visibility for all valid ($t\geq t^{i}$) times.
@@ 3.1 Transformer formulation
We implement the tracker as a transformer neural network (Figs. 2 and 3) $\Psi:G\mapsto O$. The goal of this transformer is to improve an initial estimate of the tracks. Tracks are encoded as a grid of input tokens $G^{i}_{t}$, one for each track $i=1,\dots,N$, and time $t=1,\dots,T$. The updated tracks are expressed by a corresponding grid of output tokens $O^{i}_{t}$.
@@
The appearance of the tracks is captured by feature vectors $Q_{t}^{i}\in\mathbb{R}^{d}$ (these are time-dependent to accommodate changes in the track appearance). They are initialized by sampling image features at the starting locations, and then they are updated by the neural network, as explained below.
Correlation features.
In order to facilitate matching tracks to images, we adopt the correlation features $C^{i}_{t}\in\mathbb{R}^{S}$ of RAFT [43]. Each $C^{i}_{t}$ is obtained by comparing the track features $Q_{t}^{i}$ to the image features $\phi_{s}(I_{t})$ around the current estimate $\hat{P}_{t}^{i}$ of the tracks’s location. Specifically, the vector $C_{t}^{i}$ is obtained by stacking the inner products $[C_{t}^{i}]_{s\delta}=\langle Q^{i}_{t},~{}\phi_{s}(I_{t})[\hat{P}_{t}^{i}/ks+\delta]\rangle,$ where $s=1,\dots,S$ are the feature scales and $\delta\in\mathbb{Z}^{2}$, $\|\delta\|_{\infty}\leq\Delta$ are offsets. The image features $\phi_{s}(I_{t})$ are sampled at non-integer locations by using bilinear interpolation and border padding. The dimension of $C^{i}_{t}$ is $\big{(}2\Delta+1)^{2}S=196$ for our choice $S=4;\Delta=3$.
@@
The input tokens $G(\hat{P},\hat{v},Q)$ code for position, visibility, appearance, and correlation of the tracks and is given by the following concatenation of features:
@@
Iterated transformer applications.
@@
and sets $\hat{P}^{(m+1)}=\hat{P}^{(m)}+\Delta\hat{P}$ and $Q^{(m+1)}=Q^{(m)}+\Delta Q.$ The visibility mask $\hat{v}$ is not updated by the transformer, but once after the last of the $M$ applications of the transformer as $\hat{v}^{(M)}=\sigma(WQ^{(M)})$, where $\sigma$ is the sigmoid activation function and $W$ is a learned matrix of weights. We found iterative updates for the visibility did not further improve the performance, likely due to the fact that visibility highly depends on predicting an accurate location first.
@@
For $m=0$, the estimates $\hat{P}^{(0)}$, $v^{(0)}$ and $Q^{(0)}$ are initialized by broadcasting the initial values $P^{i}_{t^{i}}$ (the location of query point), $1$ (meaning visible) and $\phi(I_{t^{i}})[P^{i}_{t^{i}}/k]$ (the appearance of the query point) to all times $t=1,\dots,T$.
Figure 4: Tracked and support points. We visualize different configurations of points along with their tracks. Points represent the start of each track, which clearly illustrates grids of points in (c) and (d). (a) A single target point in a sequence from TAP-Vid-DAVIS. (b) All target points — observe that they tend to cluster on objects of interest, potentially giving our tracker an ’unfair’ advantage compared to single-track methods. (c) A local grid of support points. (d) A global grid of support points. (e) SIFT support points. We evaluate CoTracker by combining different configurations of support points, as described in Sec. 3.5
3.2 Transformer architecture and virtual tracks

### v3

# https://co-tracker.github.io/

@@ Abstract
We introduce CoTracker, a transformer-based model that tracks a large number of 2D points in long video sequences. Differently from most existing approaches that track points independently, CoTracker tracks them jointly, accounting for their dependencies. We show that joint tracking significantly improves tracking accuracy and robustness, and allows CoTracker to track occluded points and points outside of the camera view. We also introduce several innovations for this class of trackers, including using token proxies that significantly improve memory efficiency and allow CoTracker to track 70k points jointly and simultaneously at inference on a single GPU. CoTracker is an online algorithm that operates causally on short windows. However, it is trained utilizing unrolled windows as a recurrent network, maintaining tracks for long periods of time even when points are occluded or leave the field of view. Quantitatively, CoTracker substantially outperforms prior trackers on standard point-tracking benchmarks. Code and model weights are available at https://co-tracker.github.io/
Keywords:
Point tracking Motion estimation Optical flow
Figure 1: CoTracker is a new quasi-dense point tracker. It can track 70k points jointly on a single GPU, exploiting dependencies between different tracks to enhance performance. It has exceptional long-term tracking performance even in the presence of occlusions or when points leave the camera view (examples from DAVIS [46], top row displays object tracks, bottom row shows displacements on a dense regular grid. In frame 30, CoTracker successfully tracks the driver’s head even when it exists the camera view).
@@ 1 Introduction
We are interested in estimating point correspondences in videos containing dynamic objects and a moving camera. Establishing point correspondences is an important problem in computer vision with many applications [1, 59, 33, 65]. There are two variants of this problem. In optical flow, the goal is to estimate the velocities of all points within a video frame. This estimation is performed jointly for all points, but the motion is only predicted at an infinitesimal distance. In point tracking, the goal is to estimate the motion of points over an extended period of time. For efficiency, trackers usually focus on a sparse set of points and treat them as statistically independent. This is the case even for recent techniques such as TAPIR [19] and PIPs++ [69], which employ modern architectures such as transformers and can track points even in the presence of occlusions. However, points often have strong statistical dependencies (e.g., because they belong to the same object), which offers an opportunity for improvement.
We hypothesise that accounting for the dependency between tracked points can significantly improve tracking performance. To explore this idea, we introduce CoTracker, a new tracker that supports joint estimation of a very large number of tracks simultaneously, utilizing a transformer-based architecture and attention between tracks. As we show in ablations, tracking points jointly significantly improves the tracking accuracy, especially when points are occluded. In fact, we find that we can improve tracking quality further by adding more points to the tracker than required by the user. These additional support points expand the context of the tracker, porting to point tracking the idea of using context, common in visual object tracking [32, 37, 53].
@@
In addition to joint tracking, CoTracker incorporates additional architectural design innovations. The network is a transformer operating in a sliding window fashion on a two-dimensional token representation, where dimensions are time and the set of tracked points [2]. Via suitable self-attention operators, the transformer considers each track as a whole for the duration of a window and can exchange information between tracks, thus exploiting their dependencies. However, attention can be expensive when there are many tracks. Hence, we introduce the concept of proxy tokens to point tracking that act similarly to registers [17, 30, 44] while also reducing the memory complexity. These tokens are processed as if they were a small number of additional tracks and allow the switch from expensive self-attention between tracks to efficient cross-attention between tracks and proxies. In this way, CoTracker can jointly track a near-dense set of tracks on a single GPU at inference.
@@
CoTracker is designed as an online tracker that operates on relatively short windows of frames. Within a window, the tracks are initialized with queried points, and the network is tasked with progressively refining these initial estimations through iterative applications of the transformer. The windows partially overlap and communicate, similar to a recurrent network. Each subsequent overlapping window starts with refined predictions of the previous window and updates tracks for the new frames. We optimize the recurrent application of the network via unrolled training, porting this concept from recurrent networks and visual object tracking [26, 22] to point tracking. In this way, CoTracker achieves excellent long-term tracking performance as well; in particular, joint and recurrent tracking allows points to be tracked through long occlusions.
We train CoTracker on the synthetic TAP-Vid-Kubric [18] and evaluate it on TAP-Vid-{DAVIS, RGB-Stacking}, PointOdyssey [69] and DynamicReplica [35]. Our architecture works well for tracking single points and excels for groups of points, obtaining state-of-the-art tracking performance in several benchmarks compared to prior trackers. In summary, our contributions are the following: (i) we introduce the concept of joint point tracking by sharing information between tracked points through an attention mechanism; (ii) we propose to use support points to provide additional context for point tracking; (iii) we reduce the model’s memory complexity with proxy tokens; and (iv) we propose an unrolled model training strategy, which further improves tracking and occlusion accuracy.
Figure 2: Non-joint tracking (top row) vs. joint tracking (bottom). Background points are colored cyan while foreground points are magenta. Non-joint point tracking causes background points to follow object motion (cyan points in later frames), while tracking points together results in better tracks for foreground and background points.
@@ 2 Related work
Optical flow approximates dense instantaneous motion. Originally approached by studying the brightness constancy equation [27, 40, 6, 7], starting with FlowNet [20, 29] optical flow has been tackled using deep learning. More recently, DCFlow [64] has introduced the computation of a 4D cost volume, used in most follow-up works [56, 61, 58]. Notable is RAFT [58], which introduced incremental flow updates and inspired several follow-up works [34, 63, 34, 66]. CoTracker also uses 4D cost volumes and iterated updates but is applied to tracking.
Transformers [60] have also been applied to the optical flow problem [28, 50, 67]. Flowformer [28] drew inspiration from RAFT and proposed a transformer-based approach that tokenizes the 4D cost volume. GMFlow [67] replaced the update network with a softmax with self-attention for refinement. Perceiver IO [30] proposed a unified transformer for several tasks, including optical flow.
Optical flow can be applied to tracking by integrating predictions over time, but this approach accumulates errors in the form of drift [24], which motivates developing architectures like CoTracker that track points over longer horizons.
@@
Several authors [47, 49, 31, 41] have extended optical flow to multiple frames. Early methods used Kalman filtering [13, 21] to encourage temporal consistency. Modern multi-frame methods produce dense flow. RAFT [58] can be applied in a warm-start manner [54, 57] for multi-frame optical flow estimation. However, these methods are not designed for long-term tracking and do not consider points occluded for a long time. VideoFlow [49] extends optical flow to three and five consecutive frames by integrating forward and backward motion features during flow refinement. MFT [45] conducts optical flow estimation between distant frames and chooses the most reliable chain of optical flows. Our tracker scales to produce semi-dense tracks.
Visual object tracking.
Before the emergence of deep learning, some authors proposed handcrafted joint trackers [5, 48], but few have considered doing so using deep networks like CoTracker. Our work is weakly related to multiple object tracking [11] where tracking through occlusions [68], with changes in appearance [42], and with temporal priors [51] have been extensively studied.
Some recent point trackers have been inspired by visual object tracking that predict the target states by extracting and combining features extracted from source and target image regions [3, 16, 36, 4, 15]. Transformers have also been applied in visual object tracking [12, 14]. However, our focus is the tracking of points, including background points, not objects.
@@
Particle Video [48] introduced the problem of Tracking Any Point in a video (TAP), and PIPs [24] proposed a better model which can track through occlusions more reliably by predicting tracks in a sliding window and restarting them from the last frame where the point was visible. However, PIP cannot track a point beyond the duration of a window. TAP-Vid [18] introduced a new benchmark and a simple baseline for TAP. While the original Particle Video does track points jointly, PIPs and TAP-Vid track points independently (in parallel). Inspired by TAP-Vid [18] and PIPs [24], TAPIR [19] further improved tracking with a two-stage, feed-forward point tracker that uses TAP-Vid-like matching and PIPs-like refinement.
PointOdyssey [69] addressed long-term tracking with PIPs++, a simplified version of PIPs, and introduced another benchmark for long-term tracking. However, PIPs++ still tracks points independently. OmniMotion [62] optimizes a volumetric representation for each video, refining correspondences in a canonical space. However, this approach requires test-time optimization, which, due to its cost, is not suitable for many practical applications, especially online ones.
Trackers and optical flow models are often trained using synthetic datasets [35, 43, 20, 55, 8, 69], as annotating real data can be challenging. Synthetic datasets provide accurate annotations, and training on them has demonstrated the ability to generalize to real-world data [43, 55].
@@ 3 CoTracker
Our goal is to track 2D points throughout the duration of a video $V=(I_{t})_{t=1}^{T}$, which is a sequence of $T$ RGB frames $I_{t}\in\mathbb{R}^{3\times H\times W}$. The goal of the tracker is to predict $N$ point tracks $P^{i}_{t}=(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}$, $t=t^{i},\dots,T$, $i=1,\dots,N$, where $t^{i}\in\{1,\dots,T\}$ is the time when the track starts. The tracker also predicts the visibility flag $v^{i}_{t}\in\{0,1\}$, which tells if a point is visible or occluded in a given frame. To make the task definition unambiguous [18], we assume that each point is visible at the start of the track (i.e., $v^{i}_{t_{i}}=1$). The tracker is thus given as input the video $V$ and the starting locations and times $(P^{i}_{t^{i}},t^{i})_{i=1}^{N}$ of $N$ tracks, and outputs an estimate $(\hat{P}^{i}_{t}=(\hat{x}^{i}_{t},\hat{y}^{i}_{t}),\hat{v}^{i}_{t})$ of the track locations and visibility for all valid times, i.e., $t\geq t^{i}$.
Figure 3: CoTracker architecture. We compute convolutional features $\phi(I_{t})$ for every frame and process them with sliding windows. To initialize track features $Q$, we bilinearly sample from $\phi(I_{t})$ with starting point locations $P$. Locations $P$ also serve to initialize estimated tracks $\hat{P}$. See Fig. 4 for a visualization of one sliding window.
@@ 3.1 Transformer formulation
We implement the tracker as a transformer neural network (Figs. 3 and 4) $\Psi:G\to O$. The goal of this transformer is to improve an initial estimate of the tracks. Tracks are encoded as a grid $G$ of input tokens $G^{i}_{t}$, one for each track $i=1,\dots,N$, and time $t=1,\dots,T$. The updated tracks are expressed by a corresponding grid $O$ of output tokens $O^{i}_{t}$.
@@
The appearance of the tracks is captured by feature vectors $Q_{t}^{i}\in\mathbb{R}^{d}$ (these are time-dependent to accommodate changes in the track appearance). They are initialized by broadcasting image features sampled at the starting locations and are updated by the neural network, as explained below.
Spatial correlation features.
In order to facilitate matching tracks to images, we adopt correlation features $C^{i}_{t}\in\mathbb{R}^{S}$ similar to RAFT [58]. Each $C^{i}_{t}$ is obtained by comparing the track features $Q_{t}^{i}$ to the image features $\phi_{s}(I_{t})$ around the current estimate $\hat{P}_{t}^{i}$ of the track’s location. Specifically, the vector $C_{t}^{i}$ is obtained by stacking the inner products $[C_{t}^{i}]_{s\delta}=\langle Q^{i}_{t},~{}\phi_{s}(I_{t})[\hat{P}_{t}^{i}/ks+\delta]\rangle,$ where $s=1,\dots,S$ are the feature scales and $\delta\in\mathbb{Z}^{2}$, $\|\delta\|_{\infty}\leq\Delta$ are offsets. The image features $\phi_{s}(I_{t})$ are sampled at non-integer locations by using bilinear interpolation and border padding. The dimension of $C^{i}_{t}$ is $\big{(}2\Delta+1)^{2}S=196$ for our choice $S=4;\Delta=3$.
@@
The input tokens $G(\hat{P},\hat{v},Q)$ code for position, visibility, appearance, and correlation of the tracks. They are given by the following concatenation of features with added positional encodings:
@@
Iterated transformer application.
@@
and sets $\hat{P}^{(m+1)}=\hat{P}^{(m)}+\Delta\hat{P}$ and $Q^{(m+1)}=Q^{(m)}+\Delta Q.$ The visibility mask $\hat{v}$ is not updated iteratively, but only once after the last transformer application as $\hat{v}^{(M)}=\sigma(WQ^{(M)})$, where $\sigma$ is the sigmoid activation function and $W$ is a learned matrix of weights. We found that updating the visibility flag iteratively did not improve performance, likely due to the fact that predicting visibility requires predicting an accurate location first.
@@
For $m=0$, the position, visibility and appearance estimates $\hat{P}^{(0)}$, $v^{(0)}$ and $Q^{(0)}$ are initialized by broadcasting their query point’s initial location $P^{i}_{t^{i}}$, visibility $v^{i}_{t^{i}}=1$ (meaning visible) and appearance $\phi(I_{t^{i}})[P^{i}_{t^{i}}/k]$ to all times $t=1,\dots,T$.
Figure 4: CoTracker architecture. Visualization of one sliding window with $M$ updates. During one iteration, we update point tracks $\hat{P}^{(m)}$ and track features $Q^{(m)}$. $Q^{(0)}$ is initialized with the initially sampled features $Q$ for all sliding windows, $\hat{P}^{(0)}$ with the starting locations for the first window. For other windows, $\hat{P}^{(0)}$ starts with predictions for frames processed in the preceding sliding window, and with the last predicted positions for the unseen frames. We compute visibility $\hat{v}$ after the last update $M$.
3.2 Transformer architecture and proxy tokens